### 04 Mar 2024 v2

# T-PRIME: Transformer-based Protocol Identification for Machine-learning at the Edge

Abstract
Spectrum sharing allows different protocols of the same standard (e.g., 802.11 family) or different standards (e.g., LTE and DVB) to coexist in overlapping frequency bands. As this paradigm continues to spread, wireless systems must also evolve to identify active transmitters and unauthorized waveforms in real time under intentional distortion of preambles, extremely low signal-to-noise ratios and challenging channel conditions. We overcome limitations of correlation-based preamble matching methods in such conditions through the design of T-PRIME: a Transformer-based machine learning approach. T-PRIME learns the structural design of transmitted frames through its attention mechanism, looking at sequence patterns that go beyond the preamble alone. The paper makes three contributions: First, it compares Transformer models and demonstrates their superiority over traditional methods and state-of-the-art neural networks. Second, it rigorously analyzes T-PRIME’s real-time feasibility on DeepWave’s AIR-T platform. Third, it utilizes an extensive 66 GB dataset of over-the-air (OTA) WiFi transmissions for training, which is released along with the code for community use. Results reveal nearly perfect (i.e. $>98\%$) classification accuracy under simulated scenarios, showing $100\%$ detection improvement over legacy methods in low SNR ranges, $97\%$ classification accuracy for OTA single-protocol transmissions and up to $75\%$ double-protocol classification accuracy in interference scenarios.
Index Terms:
deep learning, protocol classification, edge computing
I Introduction
The increasing demand for wireless services has caused a scarcity of spectrum resources [1]. This results in congested wireless spectrum environments as various communication protocols coexist in the same frequency bands [2]. Unauthorized transmissions further raise security concerns, posing risks to critical operations [3]. Detecting diverse protocols in crowded spectrums allows for intelligent strategies to mitigate interference, improve spectral efficiency, and enhance overall wireless system performance, benefiting regulatory bodies, network operators, and researchers.
Figure 1: T-PRIME system overview and receiver design.
$\bullet$ Problems with legacy methods of protocol classification. Legacy methods require extensive prior knowledge of protocols, leading to challenges in detecting new protocols. Even with full protocol knowledge, correlation-based methods suffer from reduced detection accuracy in low SNR conditions and when multiple protocols overlap in frequency. To address these challenges, T-PRIME (shown in Fig. 1) adopts a Machine Learning (ML) based approach, performing protocol classification without requiring specific channel boundaries or symbol rates, even with overlapping signals and low SNR conditions.
Legacy protocol detection methods are integrated into the RF receiver network interface card, enabling fast preamble correlations in hardware. However, updating the system for new protocols leads to backward compatibility issues and potential detection errors in challenging wireless channels. On the other hand, software-defined radio (SDR) systems offer flexibility but introduce higher latencies due to data transfer and complexities in I/O and buffer management. This paper demonstrates that using edge devices with CPU and GPU on a system-on-a-module (SOM), along with careful software design, can overcome these limitations and enable real-time processing for complex ML wireless applications on an SDR-based edge device.
$\bullet$ Problems with WiFi protocol classification. We showcase the potential of SDR based, ML enabled architectures by focusing on a challenging wireless scenario involving protocol classification. Specifically, we classify different 802.11 WiFi standards, including 802.11b, 802.11g, 802.11n, and 802.11ax. These standards share similar preamble structures and can operate with the same MCS, frequency bands, channel bandwidths, and transmit power. Importantly, this classification is distinct from and more difficult than prior works [1, 2, 3], which classify diverse protocols like LTE, WiFi, and Bluetooth from a candidate pool and lack a real time deployment.
$\bullet$ Proposed solution: T-PRIME. We demonstrate the effective use of Transformer neural networks [4] for the above WiFi protocol classification problem. Transformers have gained significant interest in the deep learning community for their ability to ingest long sequences of data and efficiently correlate temporal features far from each other, which is particularly useful for natural language processing tasks [5]. However, their application to wireless communications is still in a nascent stage, with some works focusing on modulation classification [6, 7, 8]. Although Transformer-based models excel at processing 1D sequential and time-dependent information, adapting them to work with 2D sequences of IQ samples is not straightforward. Additionally, it is crucial to evaluate ML models intended to be deployed as real-time classifiers on resource-constrained devices to understand the trade-offs between accuracy and delay. Our contributions are as follows:
  • We create, and publicly release [9], a protocol classifier called T-PRIME based on Transformer architecture and compare it to both classical, preamble based processing and ML based modulation detectors using a large data set of simulated WiFi signals.
  • We rigorously evaluate the performance of each of these protocol classification methods under various Single-Input-Single-Output (SISO) wireless channel settings, including randomized channel realization, different signal-to-noise ratio (SNR) levels, and multiple channel models. In all these settings, T-PRIME outperforms state-of-the-art legacy and ML solutions available for classification of raw IQ signals.
  • We implement T-PRIME on DeepWave’s Artificial Intelligence Radio Transceiver (AIR-T) and evaluate performance in realistic scenarios. Drawing from this experience, we provide insights that will allow building a generic ML-based wireless signal processing pipeline on SOM architectures that operates in near real-time using sequences of time-domain IQ samples.
  • We obtain and release the first-of-its-kind dataset comprising over-the-air (OTA) WiFi-signals for different protocol versions. This dataset includes transmissions collected in diverse environments, encompassing both single and overlapping transmission scenarios.
II Related Work
$\bullet$ Preamble correlation Methods. Frame preamble usage for protocol identification and synchronization involves detecting a predefined symbol sequence at the start of a transmitted signal. For example, the 802.11b long preamble consists of 128 scrambled 1s and 16 SFD marker bits, while the 802.11n preamble includes at least 9 symbols spread across all 52 OFDM sub-carriers.
Traditional signal processing techniques use energy detection and correlation [10, 11], requiring prior knowledge of the unique preamble sequences. However, they can be vulnerable to noise, interference, and channel variations. In Sec. IV-A, we demonstrate how legacy approaches fail at low SNR, while T-PRIME achieves highly accurate protocol classification under the same conditions.
$\bullet$ ML-based methods. Protocol classification via ML methods is still in a nascent stage. Zhang et al.  [1] perform protocol detection between WiFi, LTE, and 5G signals using a multi-layer bidirectional RNN. They use 512 IQ pairs with a sliding window as inputs to their RNN. Similarly, Schmidt et al. use a CNN to classify between WiFi, Bluetooth, and ZigBee signals [2]. Jagannath and Jagannath [3] perform both modulation classification and signal classification via CNNs by distinguishing between a wide range of different protocols including AM Radio, Bluetooth, WiFi, and IEEE 802.15.4 (Zigbee). However, these approaches do not address the challenge of classifying similar protocols, such as different WiFi standards.
[…]
III T-PRIME Design
Transformers, as detailed in [20], employ attention mechanisms, or activation masks, which filter information within a layer for transmission to the subsequent layer. In self-attention, the network generates this mask based on input and/or preceding layers. Transformers excel in modeling varied dependencies among sequence elements, independent of their positions in input or output sequences, unlike RNNs. Crucially, these adaptable dependencies in Transformers are shaped by and learned from input data.
Figure 2: Transformer-based model architecture.
The general Transformer architecture, as described by Vaswani et al. [4], is composed of several stacked transformer layers shown in Fig. 2, typically forming an encoder-decoder structure. Each layer has two sub-layers: a multi-head attention mechanism and a feed-forward network. The input data, along with positional encoding, is fed into the Transformer. Next, it is passed to the multi-head attention mechanism. There are residual connections and normalizations around the two sub-layers. Implementations generally differ in where the normalization occurs and in the specific implementation of the feed-forward layer.
Figure 3: Transmission split into sequences, slices and, tokens.
$\bullet$ Adapting Transformer for IQ Input. The classical Transformer architecture requires a token sequence as input, typically achieved in NLP tasks through word embeddings and positional information. However, feeding IQ data to the Transformer is not straightforward, leading to complex approaches in prior work [7, 8, 6]. Initially, we used spectrograms and a vision Transformer [21], but this yielded lower performance compared to a baseline CNN [12].
[…]
$\bullet$ Transformer Architecture Design. Our proposed Transformer architecture (Fig. 2) omits positional encoding as it did not affect accuracy. Since our task is classification, we utilize only the encoder part of the Transformer. Contextualized embeddings from the encoder layers flow into a Fully-Connected layer and an Output layer representing the four classes in our classification. The architecture includes an initial normalization layer, LogSoftmax as the final activation function and NLLLoss as loss function.
We evaluate T-PRIME using a systematic two-phase approach. In the first phase, we generate a large synthetic dataset to compare T-PRIME with the legacy correlation-based WiFi preamble detection algorithm and ML models from prior work. T-PRIME achieves superior results on synthetically generated datasets compared to these other approaches. In the second phase, we conduct an in-depth evaluation of T-PRIME on a more realistic dataset consisting of over-the-air transmissions collected in diverse and complex RF environments. Additionally, we implement T-PRIME in the AIR-T testbed to demonstrate its accuracy and real-time inference latency.
IV Offline Evaluation Over Synthetic Data
IV-1 Dataset Description
We used MATLAB’s WLAN Waveform Generator to generate 802.11b, 802.11g, 802.11n and 802.11ax waveforms. This dataset contains baseband IQ samples according to the specifications shown in Table I. For each protocol, we generate $2000$ Physical layer Protocol Data Units (PPDU) bursts, each comprising of multiple packets carrying a $1000$ bit random payload, saved into individual files. As 802.11b does not support a bandwidth/sampling rate of $20$MHz, the waveform is upsampled from $11$MHz to $20$MHz to ensure waveforms have the same channel bandwidth and ensure the problem is challenging for the classifier.
Standard Bandwidth MCS Modulation #PPDU/burst #IQ/burst
802.11b* 11MHz QPSK DSSS 1 18112
802.11g 20MHz 16-QAM (½ & ¾) OFDM 4 & 4 32960
802.11n 20MHz 16-QAM (½ & ¾) OFDM 5 & 6 31340
802.11ax 20MHz 16-QAM (½ & ¾) OFDMA 5 & 6 31640
TABLE I: Synthetic dataset specifications. 802.11b* is upsampled from 11MHz to 20MHz for consistency with other signals. PPDUs per burst are reported for each coding rate.
IV-2 Training Procedure
Once the burst signals have been generated, we use a PyTorch data generator to train our models that operates on a per-training batch basis to dynamically load into memory each file individually and simulate transmission through a channel model at an SNR level randomly chosen. We consider the following options for (a) channel models: TGn, TGax, Rayleigh, no channel impairment; and (b) a random SNR level defined between $-30$ dB and $30$ dB, applied using Additive White Gaussian Noise (AWGN). We develop a Python wrapper for Matlab’s channel models standard compliant implementation and employ Model-B multi-path delay profile for TGn and TGax with device distance of $3$m, while for Rayleigh model we use a single Non-Line Of Sight (NLOS) path with an average path gain of $-3$ dB and delay of $1.5$e-$9$ seconds.
Input sequences are pre-generated before training procedure for the complete set of baseband signals in the dataset. Once sequence indexes are generated for each dataset burst, they are randomly shuffled and then separated into training and validation set with a split of $80$% and $20$%. When a mini-batch of sequences is generated at training time, the complete burst signal is retrieved and passed through a random channel model and AWGN. Finally, the input sequence is retrieved from the complete distorted signal and divided into slices as described in Fig. 3. Multi-path distortions thus impact the complete set of IQ samples in a batch on a per-sequence basis, simulating realistic transmission.
While dynamic signal augmentation is performed for both train and validation data, to evaluate the performance of our models in similar conditions we use an additional static test set, composed of $500$ bursts for each protocol following the configuration in Table I. Such test transmissions are pre-processed by creating multiple copies of the bursts that are passed through a random channel instance for all the considered channel models, for a total of an additional $2000$ test packet bursts. Then, at test time, we create multiple sequences from each signal copy and evaluate them on discrete SNR levels ranging between $-30$ and $30$ dB in increments of $5$ dB.
The proposed T-PRIME design was finalized using the same training procedure above and a hyperparameters tuning phase based on a grid search on the validation set. The chosen training configuration for transformers architectures is $5$ training epochs, batch size of $122$ and learning rate $2$e44-4- 4. We found that a stack of $2$ transformer layers is optimal, along with two distinct sets of values for sequence length and slice length that provided nearly perfect results. The first model, named Small (SM) Transformer, utilized an input sequence of $M=24$ and tokens with a length of $S=64$. The second, Large (LG) model, employed a sequence of $M=64$ tokens with a length of $S=128$ each. The resulting number of parameters of SM and LG architectures are $1.6$M and $6.8$M, respectively.
IV-A Comparison with state-of-the-art legacy and ML methods
We first show how T-PRIME approach can effectively improve the state-of-the-art over legacy approaches. Specifically, we use the same static test set described in Sec. IV-1 for a simulation study that processes each test transmission using the correlation-based procedure specified in the standard [10]. We employ a set of multiple decoders provided by Matlab that use the legacy method to detect the following preambles formats: 1) VHT (i.e. 802.11n), 2) HE-SU (i.e. 802.11ax), 3) Non-HT (i.e. 802.11g/b). We then process our test transmissions through these functions for each channel model considered. For each test signal, we simulate $100$ transmissions over random channel instances for each channel model and SNR levels considered in our evaluation procedure described before. To report the accuracy of legacy procedures, we consider the inverse Packet-Detection-Rate (i.e. $1$ PDR), defined as the ratio of correct preamble format detections over the total number of simulated packets transmissions, and compare it with accuracy of proposed architectures under the same statistical channel and SNR conditions. We evince from Fig. 3(a) that both our T-PRIME models drastically outperform the legacy detection mechanism, achieving near-perfect detection accuracy already at SNR $=0$ dB, while traditional methods completely fail to operate under such circumstances. This demonstrates how these models are still able to extract and learn powerful time-domain dependencies from received signals even in presence of strong noise or low reception power.
After establishing superiority of our approach over legacy methods, we compare our model to other state-of-the-art ML methods for wireless signal classification. While our problem formulation differs from modulation classification, it serves as a relevant example of architectures processing sequences of raw IQ samples for similar tasks. We compare T-PRIME against three state-of-the-art models in this domain: ResNet [15], AMCNet [18], and MCformer [8], using the hyperparameters reported as optimal by their respective authors; we also present supplementary findings obtained through additional hyperparameter exploration in Appendix. For further comparison, we trained a 1D CNN model inspired by [12] with a wider input size and slightly larger set of parameters. We used a common training configuration for all models with a batch size of 512, $M=1$ for input generation, CrossEntropyLoss as the loss function, and a learning rate of $0.001$ with exponential decay to $0.0001$ when a plateau in the validation loss is reached. Table II lists all the models we compared, along with their input size and total number of parameters.
Figure 4: (a) Comparison between SM, LG Transformer-based architectures and state-of-the-art signal classification models tested on all channel models and different SNR conditions. The LG Transformer achieves the best overall accuracy and it retains high accuracy (i.e. $>98\%$) for SNR as low as -10 dB. All models improve under optimal hyperparameter search (see Appendix), but the LG Transformer still outperforms all other methods. (b) Performance comparison for each individual channel model tested during simulation.
Model Input size ($M\times S$) # Parameters
T-PRIME LG $64\times 128$ 6.8M
T-PRIME SM $24\times 64$ 1.6M
1D CNN [12] $1\times 512$ 4.1M
ResNet [15] $1\times 1024$ 162K
AMCNet [18] $1\times 128$ 462K
MCformer [8] $1\times 128$ 78K
TABLE II: Comparison between proposed architectures and state-of-the-art MCS classification architectures.
Fig. 4 illustrates that both T-PRIME LG and T-PRIME SM consistently outperform all other methods across the entire range of evaluated signal-to-noise ratios (SNRs). Notably, T-PRIME LG achieves a remarkable $>60$% improvement over all other models at SNR $=-10$ dB, achieving 98% accuracy. From these experiments, two key points can be deduced: 1) T-PRIME’s performance benefits from a larger parameter space, following the trend of Large-Language-Models (LLMs) that use substantial parameter capacity to learn complex relationships [22]; 2) Our problem significantly differs from MCS classification, as the results suggest that larger amounts of temporal sequences are required to accurately identify the complex and less localized patterns in real-world wireless protocols, especially when protocols share the same MCS configurations. Finally, hyperparameter search improves all models (see Appendix); while, in this case, AMCNet surpasses the SM Transformer for low SNR, nevertheless the LG Transformer still consistently outperforms all competitors.
V Real-Time Implementation on AIR-T SDR
SDR-based systems offer flexibility in the receiver chain and seamless integration with AI/ML modules. However, they can introduce higher latency due to challenges like I/O and buffer management, software dependencies, and distribution of computing resources for signal processing tasks. To explore these systems’ challenges and showcase T-PRIME’s real-time efficiency, we implemented it on Deepwave Digital’s AIR-T model AIR7101, as shown in Fig 5. The AIR-T, built on NVIDIA’s Jetson TX2 module, includes a CPU with $4$ ARM A-57 cores, $2$ ARM Denver2 cores, and a Pascal $256$-core GPU with $8$ GB of shared memory [23]. This design resolves traditional SDR system issues by integrating resources on one platform, including SDR with FPGA, networking, periphery connections, an Ubuntu-based OS, and open source APIs.
Figure 5: T-PRIME running real-time classification in an AIR-T receiver. 802.11g protocol is shown to be correctly detected in this instance.
For real-time protocol classification, low latency is crucial. Hence, we designed a pipeline (Fig. 6) with three primary functions running in parallel: receiver, signal processing, and ML inference. This parallelization removes concurrent dependencies and reduces prediction delay. It also allows individual function modification or improvement without affecting others. For instance, we deployed the same Transformer model using two frameworks (PyTorch and TensorRT) transparently. Each thread includes timing functions to calculate the average time for function completion. Additionally, we established two thread-safe FIFO queues, q1 between receiver and signal processing threads, and q2 between signal processing and ML inference threads. Below, we briefly describe each of these three threads; see also our code [9] for further details.
Figure 6: T-PRIME pipeline overview showing three steps of the receiver, signal processing and ML processing modular blocks.
V-1 Receiver Thread
The main function of the receiver thread is to capture the raw OTA signals. The receiver thread defines several parameters such as sample rate ($31.25$ MHz), gain mode, and frequency for the AIR-T receiver using SoapySDR APIs. After initializing the SDR, it creates a data buffer to receive samples from the SDR. After these one time initialization steps are complete, the receive thread enters a continuous loop that reads the buffer, ensures it received the expected number of samples, and writes the vector of samples to q1 if there is space available.
V-2 Signal Processing Thread
The signal processing thread’s main function is to down-sample the received signal to $20$ MHz and convert it to the format required by the ML inference thread. After reading the raw sample vector from q1, the thread converts the data type from interleaved shorts to numpy.complex64, normalized to the range of $[-1,1]$. Then, we down-sample the signal from $31.25$ MHz to $20$ MHz, the standard channel bandwidth used in 802.11 WiFi protocols. Finally, the complex vector is converted into a real vector by interleaving the I and Q (real and imaginary) components, as shown in Fig. 3. If space is available, the thread writes this vector to q2.
V-3 ML Inference Thread
This thread is the core of our pipeline, performing the actual ML inference on the input samples. We deployed two different versions of this thread; a PyTorch version and a TensorRT version. TensorRT [24] uses techniques such as quantization, layer and tensor fusion, kernel tuning to optimize DL inference. Regardless of the optimization used, the thread follows the same process. After loading the model, it enters a continuous loop. If there is data available in q2 it reads that into a tensor to use as the model input, classifies, and appends the classification label to a text file.
To handle varying signal power caused by propagation effects, we incorporate a power normalization block in the ML inference thread. This function normalizes the total power of received signals to a nominal power of 1W. Given a generic discrete complex signal $\mathbf{s}\in\mathbb{C}^{N}$ of $N$ samples, we can obtain its normalized version $\mathbf{\hat{s}}$ as follows: $\mathbf{\hat{s}}={\mathbf{s}}/{\sqrt{\frac{\sum_{i=1}^{N}\left|s_{i}\right|}{N}}}$ where the denominator is the Root-Mean-Square (RMS) value of the incoming signal. This block makes our model more robust to power and SNR variations, eliminating the need to collect data samples and train our models under multiple power profiles.
VI T-PRIME AIR-T Testbed Evaluation
We evaluated the performance of our T-Prime implementation through: (a) the creation of an Over-the-Air (OTA) dataset, collected from the deployment of the testbed in five different locations, and use this dataset to conduct additional offline evaluation of T-Prime, (b) power and latency profile characterization of our implementation, (c) real-time, live tests of a deployment of our entire pipeline evaluating our models in a new environment. We describe each of these experiments below.
VI-A Experimental Setup
VI-A1 Over-the-Air (OTA) Experiments
In order to evaluate the performance of our proposed approach in a real-world setting, we collect OTA transmissions in a variety of conditions, including different indoor scenarios and transmission power levels. To collect this data, we re-use the same receiver and signal processing blocks. Instead of passing the resulting vectors to the ML inference block, we write them to a file for offline training and testing.
$\bullet$ Dataset collection. We collected a large OTA dataset, providing a robust collection of real-world wireless channel conditions. Table III summarizes the environments for each collection set. Room A is a large open classroom, while rooms B and C are small lab rooms. The level of interference was assessed based on visible WiFi access points and other known transmitters in the $2.4$ GHz band. The level of scatter considers the ratio of room dimensions to the space used by equipment and furniture. Each room’s data was collected on two separate days (_1 vs _2). We also collected noise/background samples to measure the noise floor and test classifier performance when no active transmission occurred.
Dataset Distance Scatter Interference #Captures per protocol
ax b n g noise
RM_A_1 3m Low Low 200 206 200 200 0
RM_A_2 2m Low Medium 200 200 200 200 0
RM_B_1 1m Med Medium 73 200 200 200 0
RM_B_2 1m Med Medium 500 500 500 500 500
RM_C_1 1m High High 100 100 100 100 0
RM_C_2 1m High High 500 500 500 500 100
[…]
$\bullet$ Single-protocol OTA dataset. The above data collection results in an offline OTA dataset summarized in Table III. We split this into a training and test datasets in two ways. In our time-split, for each day of data collection, we consider the first $80$% of samples collected per protocol as a training set, and the remaining $20$% as a test set. In our scenario-split, we perform a cross-validation study across all the location-specific datasets. We train our classifier on all OTA collections except one, which we use in its entirety to test the resulting model. In both cases the test data channel conditions are unseen during the training phase.
$\bullet$ Overlapping-protocol OTA datasets. We stress-tested T-PRIME in the presence of spectral overlaps, a scenario where traditional preamble detection techniques struggle due to frame collisions corrupting the preamble structure. In RM_C, we collected two additional datasets with signals transmitted concurrently, overlapping at $25\%$ and $50\%$ spectral ratios. We considered six different overlapping protocol configurations (see Table IV). The datasets differed in the observed wireless spectrum resolution, achieved by varying the receiver sampling rate and central frequency.
Configuration Incumbent Interferer
C1 802.11g 802.11n
C2 802.11b 802.11ax
C3 802.11b 802.11g
C4 802.11b 802.11n
C5 802.11g 802.11ax
C6 802.11n 802.11ax
TABLE IV: Protocols configuration for overlapping cases.
Our goal was to test if our Transformer classifier could recognize both incumbent signals and the protocol of the interfering transmission using raw IQ time-domain samples. We defined two additional datasets with different receiver configurations: O1) the incumbent and receiver tuned to the same WiFi channel (WiFi Ch. 7) and sharing the same bandwidth, and O2) the receiver observing a wider bandwidth and tuned to an arbitrary frequency ($2.45$ GHz) to observe both overlapping signals in the frequency domain. Table V summarizes the parameters used for the overlapping data collections, and Fig. 7 visualizes each receiver configuration with C5 type (802.11g / 802.11ax) transmissions as an example.
Dataset $R$ $f_{s}^{\text{Rx}}$ $f_{c}^{\text{Rx}}$ $f_{c}^{\text{Tx1}}$ $f_{c}^{\text{Tx2}}$Capt./Config #IQ/Capt.
RM_C_O1 25% 20 2.442 2.442 2.457 200 198080
RM_C_O1 50% 20 2.442 2.442 2.452 200 198080
RM_C_O2 25% 62.5 2.45 2.442 2.457 200 309500
RM_C_O2 50% 62.5 2.45 2.442 2.452 200 309500
TABLE V: OTA overlapping dataset parameters. $f_{s}^{\text{Rx}}$ (expressed in MHz) is receiver sampling rate. $f_{c}^{\text{Rx}}$, $f_{c}^{\text{Tx1}}$ and $f_{c}^{\text{Tx2}}$ (expressed in GHz) are receiver central frequency, central frequencies of first transmitter (incumbent) and central frequency of second transmitter, respectively.
(a) $f_{s}^{\text{Rx}}=20$, $R=25\%$
(b) $f_{s}^{\text{Rx}}=20$, $R=50\%$
(c) $f_{s}^{\text{Rx}}=62.5$, $R=25\%$
(d) $f_{s}^{\text{Rx}}=62.5$, $R=50\%$
Figure 7: Example spectrograms of recorded transmissions used in the interference dataset. Samples are shown for each configurations of overlapping rate $R$ and sampling rate $f_{s}^{\text{Rx}}$ expressed in MHz.
We adapted our classifier from a single to multi-label output by replacing the last LogSoftMax activation function with Sigmoid at each output neuron and using BCELoss as loss function. We trained our classifier on all collected OTA transmissions, including single protocol and overlapping ones, using the time-split method with the same $80$/$20$ split for training/testing. Note that now we include both non-overlapping and overlapping cases, and signals with different sampling rates (e.g. _O2 datasets) in the training set. This challenging dataset forces the Transformer to learn different patterns and ultimately classify the protocols irrespective of their sampling rate.
VI-A2 Latency and Power Profile Experiments
The AIR-T has $6$ power profiles, providing several different combinations of CPUs enabled and various clock rates for the CPUs and GPU. We performed a thorough analysis of all $6$ configurations to understand the constraints of our system along with the benefit and cost of each option. We used NVIDIA’s JTOP utility to monitor power consumed while T-PRIME performed real-time protocol classification.
VI-A3 Real-Time Experiments
We also evaluate the accuracy of T-PRIME LG in classifying OTA signals in real time. For this experiment we train T-PRIME using the entire single protocol OTA dataset (Table III). We deploy T-PRIME in two new locations, never seen during training. In both rooms, the transmitter is approximately $3$m from the T-PRIME receiver. We sequentially generate and transmit each of the four 802.11 protocols in the same manner as described in Sec. VI-A1. While continuously transmitting a given protocol, we capture $1001$ live predictions made by T-PRIME. Each experiment lasts approximately $10$s.
VI-B Results
VI-B1 OTA Dataset Evaluation
First, we show the performance of our classifier on OTA signals.
$\bullet$ Single-protocol results. Fig. 8 shows that both LG and SM models achieve an average classification accuracy $\geq 99\%$ on time-split test data, with SM architecture slightly out-performing the LG one, possibly due to overfitting for the LG model. Intuitively, a smaller variation of input patterns may be easier to learn compared to randomly simulated scenarios in the synthetic dataset. Fig. 9 shows the effect of the cross-validation on scenario-split data. This highlights that RM_C constitutes a more unique and challenging environment for our protocol classifier compared to other locations.
(a) LG model
(b) SM model
Figure 8: WiFi protocol classification performance on the complete OTA test dataset for single transmissions. Comparison between Large (LG) and Small (SM) Transformer models.
(a) RM_A_1
(b) RM_A_2
(c) RM_B_1
(d) RM_B_2
(e) RM_C_1
(f) RM_C_2
Figure 9: Cross-validation using LG classifier over the different OTA data collection instances. Each plot indicates what dataset is missing from the training set and the resulting model performance on that specific dataset at inference time.
(a) W/O Power norm.
(b) W/ Power norm.
Figure 10: Impact of power normalization for low SNR conditions, i.e. $P_{\text{Tx}}=0$ dBm, on test data collected in RM_C.
$\bullet$ Overlapping-protocol results. Table VI shows the average test accuracy on the complete OTA test dataset and, separately, a detailed view of accuracy for each individual overlapping transmission dataset. We use three different metrics to determine the performance: 1) percent of “exact” classifications, i.e. correctly classifying both protocols present in the received signals, 2) percent of times we correctly recognize a “single” protocol and 3) the Area-Under-the-Curve (AUC) of the trained classifier. These results show that our model tested on RM_C_O2 is still able to correctly detect at least one protocol $94.4\%$ and $93.1\%$ of times for $25\%$ and $50\%$ cases, respectively, and even able to detect both signals at the same time with $75.7\%$ and $66\%$ of accuracy for $R=25\%$ and $R=50\%$, respectively, which is not possible using legacy wireless hardware. Finally, we note that single-protocol accuracy is nearly perfect for RM_C_O1 (i.e. $99.9\%$ in both overlap configurations), which suggests that incumbent signals that occupy most of the observed spectrum are correctly classified the majority of times.
Dataset
OTA (all) RM_C_O1 RM_C_O2
Metric $R=25\%$ $R=50\%$ $R=25\%$ $R=50\%$
% Exact (2/2) 87.5 55.8 72 75.7 66
% Single (1/2) 98.2 99.9 99.9 94.4 93.1
AUC 0.98 0.878 0.954 0.963 0.929
TABLE VI: Average test accuracy and AUC. Classifier is trained and tested over data covering all OTA locations, including single and overlap cases.
Table VII presents Precision [^1] Precision = TP/(TP+FP); TP = True Positive; FP = False Positives., Recall [^2] Recall = TP/(TP+FN); FN = False Negative. and Support samples for each individual protocol in all receiver configurations. Precision in both RM_C_O1 cases shows higher scores for protocols 802.11b, 802.11g, and 802.11n, with a particular emphasis on 802.11b and 802.11g, as these protocols have more samples as incumbents than the others. Overall, the average Precision and Recall scores for all overlapping configurations and protocols are approximately $0.87$ and $0.86$, respectively. These results demonstrate the accuracy and completeness of positive predictions achieved by the proposed multi-protocol classifier architecture in this challenging interference scenario.
Protocol Precision Recall Support
OTA (all)
802.11ax 0.94 0.93 25482
802.11b 0.94 0.98 26154
802.11n 0.9 0.88 26106
802.11g 0.91 0.91 26106
RM_C_O1 $R=25\%$ $R=50\%$ $R=25\%$ $R=50\%$ $R=25\%$ $R=50\%$
802.11ax 0.72 0.83 0.75 0.86 2808 2808
802.11b 1 1 1 0.99 2808 2808
802.11n 0.77 0.85 0.62 0.83 2808 2808
802.11g 0.86 0.96 0.78 0.83 2808 2808
RM_C_O2 $R=25\%$ $R=50\%$ $R=25\%$ $R=50\%$ $R=25\%$ $R=50\%$
802.11ax 0.98 0.97 0.97 0.9 4329 4329
802.11b 0.82 0.87 0.93 0.96 4329 4329
802.11n 0.91 0.78 0.8 0.86 4329 4329
802.11g 0.86 0.74 0.92 0.8 4329 4329
TABLE VII: Test accuracy for individual protocols in single and overlapping transmissions. Classifier is trained and tested over data covering all OTA locations, including single and overlap cases.
$\bullet$ Impact of power normalization. We also study the impact of power normalization on an additional set of data collected under different transmission powers. Specifically, for this experiment we train our model on data collected from all locations (see Table III) but we test it on a completely different set of samples collected in RM_C with transmissions performed at power $P_{\text{Tx}}=\{30,20,10,0\}$ dBm respectively. Table VIII shows the average accuracy for each of the power profiles when power normalization is applied. While the average accuracy is consistently improved for all protocols, the most significant boost in performance is on the lower SNR end, reaching a staggering $40.3\%$ performance improvement at $P_{\text{Tx}}=0$ dBm without the need to include any additional power variation samples in the dataset.
$P_{\text{Tx}}$ (dBm) W/O Power Norm. W/ Power Norm.
0 52.5 92.8
10 93.8 95.9
20 93.9 95.3
30 97.3 97.6
TABLE VIII: Average test accuracy (%) on multiple power profiles with and without signal power normalization.
VI-B2 Real-Time Latency and Power Results
[…]
0-MAXN 1-MAXQ 2-MAXP 3-MAXP 5-MAXN
ALL ALL ARM ARM
PyTorch
[…]
CPU 1551 704 807 1457 1472
DDR 1595 842 900 999 1007
GPU 1074 486 366 434 435
SOC 1000 625 625 627 628
ALL 6933 4283 4338 5195 5242
[…]
Reciever 0.4105 0.7252 0.44307 0.4103 0.4105
Signal Pro. 9.843 15.39 14.68 10.49 10.46
ML Thread 18.98 29.18 29.58 22.28 22.27
TensorRT
[…]
CPU 1579 672 825 1515 1540
DDR 904 780 958 1131 997
GPU 404 264 436 629 578
SOC 610 598 617 668 628
ALL 5170 3923 4505 5668 5426
[…]
Reciever 0.5095 0.6162 0.524 0.4797 0.5054
Signal Pro. 10.05 15.45 13.79 9.308 9.62
ML Thread 10.91 15.91 14.82 12.46 10.86
ML $*{}^{*}$ 1.96 2.51 2.09 2.11 2.12
TABLE IX: T-PRIME LG pipeline performance for each block showing the lowest latency and lowest power using different power profiles and ML inference optimization. The last line reports average inference times of T-PRIME LG for individual inference operations within the ML Thread.
$\bullet$ Receiver thread: The data buffer in the receiver thread presented a crucial I/O constraint that required careful engineering. A large buffer ensured enough samples for our model, but caused frequent buffer overflows, adding over 1s to reset the connection. On the other hand, a small buffer could lead to insufficient data for the ML model, requiring multiple reads of the buffer. Thus, we designed the buffer to hold a specific number of complex samples based on the ML model used. For the LG Transformer, we configured the buffer to capture $12,900$ complex samples.
$\bullet$ Signal processing thread: Understanding the effects of the low pass filter and re-sampling is crucial for this block. To down-sample from $31.25$MHz to $20$MHz, we perform rational re-sampling with a factor of $\frac{16}{25}$. For our LG Transformer trained with a $20$MHz bandwidth, we need $8,192$ complex samples, so the input size should be $8008,192\times\frac{25}{16}=12,$ 800. However, in practice, the output of this thread is slightly less than the required 8,192 samples. After empirical testing, we found that providing 12,900 complex samples as input gives the correct output size. This highlights both the complexity of deployed systems and the necessity for variable length inputs to the ML model.
$\bullet$ ML inference thread: The key optimization for this thread was moving from PyTorch to TensorRT. The vast performance gains in the ML thread for TensorRT model shown in Table IX, without any loss of accuracy, highlight the importance of using well designed functional blocks which enable optimization of one function without making any changes to the others.
$\bullet$ Queue length management for real-time inference:The above pipeline offers efficient use of computing resources, flexibility in optimizing each function, and reduced latency. However, it also presents potential bottlenecks with I/O and buffer management. To maintain low end-to-end latency, optimizing the queue length between blocks is crucial. Long queues guarantee input availability for each function, maximizing parallelization gains and prediction rate. However, it significantly increases latency from signal reception to prediction. On the other hand, very short queues may lead to idle time, reducing parallelization efficiency and prediction rate. Our pipeline’s progressive processing times allow us to set short queues without starving any processes. However, jitters in processing time, IO blocks, and hardware limitations occasionally impact processing times dramatically. We found that setting both q1 and q2 to a length of $2$ minimized end-to-end delay while ensuring no thread was starved.
VI-B3 Real-Time Deployment Results
(a) RM_D.
(b) RM_E.
Figure 11: T-PRIME real-time prediction results in two new environments, RM_D and RM_E, never seen during training.
Fig. 11 shows the confusion matrix for our Real-Time evaluation where we achieve an incredible $96.7$% overall accuracy. From the confusion matrix, we see lowest accuracy is for the noise. However, given that other WiFi transmitters are operating in the area, some of the incorrect classifications for noise may in fact be correctly identifying other transmissions. Regardless, it is clear that T-PRIME has excellent accuracy in real-time predictions, even in environments it has never been exposed to during training.
VI-B4 Discussion and Insights from Real-time Inference
Our implementation and testing of T-PRIME on the AIR-T device for real-time protocol classification have yielded valuable insights applicable to other AI/ML-enabled edge device classification tasks. Utilizing integrated SOM edge devices significantly reduces latencies and complexities often encountered in SDR-based systems. Parallelization plays a crucial role in increasing the prediction rate. Designing key functions as separate threads allows for more efficient resource utilization and ease of optimization for each function. Understanding I/O points and buffer management is essential for real-time systems. The ML Inference thread is the most resource-intensive, and optimizing this task offers significant performance gains.
VII Conclusion and future work
T-PRIME achieves near-perfect classification accuracy for $4$ different WiFi protocols, outperforming legacy methods in challenging conditions. T-PRIME proves superior to existing state of the art ML models at protocol detection, especially with low SNR and overlapping signals. While our current architecture employs fixed-length inputs, Transformers excel with variable length inputs, as shown by successful NLP models like ChatGPT. In the future, we will adapt our ML block to accommodate variable length inputs. As part of our real-time implementation, we identified different bottlenecks and offered insights for solutions, which can benefit other AI/ML-based classification systems running on resource-constrained edge devices. Finally, we provide public access to our code and datasets at [9].
VIII Acknowledgment
This work was funded by the Defence Science and Technology Laboratory (DSTL), a part of the UK Ministry of Defence.
Appendix
VIII-A Hyperparameter Exploration on Baseline Models
In this section, baseline models are further explored to find optimal hyperparameters in order to compare with T-PRIME. In all tables, metrics are accuracy and cross-entropy loss respectively. In Tables X, XI, XII, and XIII, we experimented with different values of slice length, batch size, and learning rate to find the optimal configuration over the synthetic dataset. In Table XIV, optimal hyperparameters for each baseline model is shown. Except ResNet, increasing slice length provided a better validation for the models. Additionally, decreasing the batch size allow the models train better, and smaller learning rates provide a better convergence. Afterwards, including T-PRIME LG and T-PRIME SM, all models are run for 50 epochs. Overall training and validation results including the fine-tuned baseline models and T-PRIME models are shown in Table XV, where M is the sequence length, and S is the slice length. The only difference between default and fine-tuned results over T-PRIME models are the number of epochs changing from 5 to 50, while the other (already optimized) hyperparameters remain same. We put the finalized comparisons that also include the runtimes in Table XVI.
Slice Length Batch Size Learning Rate Acc. (%) Loss
128 512 1e-3 68.8 0.672
512 512 1e-3 78.8 0.464
1024 512 1e-3 80.4 0.438
2048 512 1e-3 80.0 0.432
4096 512 1e-3 75.0 0.706
8192 32 1e-3 70.2 0.713
8192 64 1e-3 75.3 0.568
8192 128 1e-3 79.9 0.450
8192 256 1e-3 81.0 0.433
8192 512 1e-4 85.9 0.324
8192 512 2e-4 86.2 0.311
8192 512 5e-4 86.4 0.315
8192 512 8e-4 84.0 0.367
8192 512 1e-3 82.6 0.430
8192 512 5e-3 51.4 1.103
8192 512 1e-2 30.0 1.357
TABLE X: AMCNet Hyperparameter Exploration
Slice Length Batch Size Learning Rate Acc. (%) Loss
512 512 1e-3 66.3 0.770
1024 32 1e-3 74.5 0.561
1024 64 1e-3 77.1 0.517
1024 128 1e-4 65.3 0.786
1024 128 2e-4 73.3 0.580
1024 128 5e-4 76.6 0.516
1024 128 1e-3 78.3 0.486
1024 128 5e-3 29.2 1.362
1024 128 1e-2 27.8 1.361
1024 256 1e-3 74.6 0.550
1024 512 1e-3 71.2 0.620
1024 1024 1e-3 68.2 0.771
1024 2048 1e-3 64.2 0.814
2048 512 1e-3 65.1 0.760
4096 512 1e-3 62.8 0.814
8192 512 1e-3 67.4 0.713
12288 512 1e-3 48.2 1.060
16384 512 1e-3 34.9 1.290
TABLE XI: ResNet Hyperparameter Exploration
Slice Length Batch Size Learning Rate Acc. (%) Loss
512 512 1e-3 45.6 1.094
1024 512 1e-3 52.3 0.997
2048 512 1e-3 39.3 1.254
4096 512 1e-3 44.6 1.152
8192 4 1e-3 55.8 0.863
8192 8 1e-3 57.2 0.850
8192 16 1e-3 57.6 0.837
8192 32 1e-3 57.6 0.839
8192 64 1e-3 57.8 0.835
8192 128 1e-4 60.8 0.828
8192 128 2e-4 60.8 0.815
8192 128 5e-4 59.8 0.832
8192 128 1e-3 58.1 0.835
8192 128 5e-3 29.8 1.345
8192 128 1e-2 29.8 1.345
8192 256 1e-3 57.6 0.864
8192 512 1e-3 55.0 0.891
8192 1024 1e-3 57.0 0.942
8192 2048 1e-3 54.5 1.057
12288 512 1e-3 33.4 1.287
16384 512 1e-3 39.4 1.305
TABLE XII: 1D CNN Hyperparameter Exploration
Slice Length Batch Size Learning Rate Acc. (%) Loss
1024 16 1e-3 52.6 0.964
1024 32 1e-3 51.7 0.968
1024 64 1e-3 51.6 0.970
1024 128 1e-3 51.4 0.974
1024 256 1e-3 51.9 0.959
2048 16 1e-3 54.2 0.903
2048 32 1e-3 53.5 0.903
2048 64 1e-4 54.3 0.909
2048 64 2e-4 55.2 0.889
2048 64 5e-4 54.7 0.899
2048 64 1e-3 54.9 0.889
2048 64 5e-3 28.3 1.358
2048 64 1e-2 28.3 1.358
4096 16 1e-3 53.7 0.895
TABLE XIII: MCformer Hyperparameter Exploration
Models Slice Length Batch Size Learning Rate
Default Fine-tuned Default Fine-tuned Default Fine-tuned
AMCNet 128 8192 512 512 1e-3 2e-4
ResNet 1024 1024 512 128 1e-3 1e-3
1D CNN 512 8192 512 128 1e-3 2e-4
MCformer 128 2048 512 64 1e-3 2e-4
TABLE XIV: Default and Fine-tuned Hyperparameters of Baseline Models
Models Training Validation
Default Fine-tuned Default Fine-tuned
Acc. (%) Loss Acc. (%) Loss Acc. (%) Loss Acc. (%) Loss
T-PRIME LG 92.4 0.213 96.4 0.179 93.5 0.166 97.5 0.068
T-PRIME SM 80.6 0.448 84.0 0.405 81.4 0.421 85.3 0.334
AMCNet 68.6 0.653 93.4 0.155 68.8 0.672 93.5 0.161
ResNet 64.7 0.667 79.9 0.403 66.3 0.770 80.3 0.439
1D CNN 53.3 0.886 60.7 0.700 45.6 1.094 61.8 0.757
MCformer 48.7 1.034 56.6 0.771 50.1 0.992 57.1 0.838
TABLE XV: Overall Training and Validation Results of T-PRIME Models, and Default and Fine-tuned Baseline Models
Models Input Size (M $\times$ S) # Parameters Runtime
Default Fine-tuned Default Fine-tuned Default Fine-tuned
T-PRIME LG 64 $\times$ 128 64 $\times$ 128 6.8M 6.8M 0d 0h 28m 0d 3h 43m
T-PRIME SM 24 $\times$ 64 24 $\times$ 64 1.6M 1.6M 0d 2h 26m 0d 19h 2m
AMCNet 1 $\times$ 128 1 $\times$ 8192 462K 269M 3d 20h 37m 0d 12h 34m
ResNet 1 $\times$ 1024 1 $\times$ 1024 162K 162K 0d 7h 27m 2d 2h 48m
1D CNN 1 $\times$ 512 1 $\times$ 8192 4.1M 67M 0d 7h 25m 0d 5h 37m
MCformer 1 $\times$ 128 1 $\times$ 2048 78K 72K 2d 9h 13m 1d 11h 3m
TABLE XVI: Comparisons between T-PRIME Models, and Default and Fine-tuned Baseline Models
VIII-B Test Results on Channel Models
In this section, T-PRIME models and optimized baseline models are compared on each individual channel model and all channel models over different values of SNR. Fig. 12 illustrates that even after the hyperparameter exploration, proposed T-PRIME LG model with fewer parameters in comparison, still consistently outperforms other optimized baseline models, achieving $96.7\%$ accuracy at SNR = $-15$ dB. Despite being outperformed by AMCNet and 1D CNN on several SNR conditions, T-PRIME SM still performs well with very few parameters compared to huge amount of parameters of these baseline models.
Figure 12: (a) Comparison between SM, LG Transformer-based architectures and fine tuned baseline models tested on all channel models and different SNR conditions. The LG Transformer achieves the best overall accuracy and it retains high accuracy (i.e. $>96.7\%$) for SNR as low as -15 dB. (b) Performance comparison for each individual channel model tested during simulation.
VIII-C Detailed accuracy metrics for multi-class classification on overlapping datasets
In this section, we present a detailed analysis of the classification accuracy of T-PRIME LG across various overlapping transmission configurations outlined in Table IV. We introduce three distinct metrics for these results, one of which is innovative compared to the metrics presented in Table VI.
We emphasize two specific accuracy metrics: Exact Accuracy and Single Accuracy. Exact Accuracy pertains to precisely identifying all and only the protocols being transmitted, while single accuracy denotes the percentage of transmissions where we successfully detect at least one of the transmitted protocols, regardless of any incorrect detection of protocols not present in the signal.
The new metric, referred to as Single Exact Accuracy, quantifies instances in which we detect at least one of the transmitted protocols while simultaneously avoiding the detection of any non-transmitted protocols. Specifically, this metric recognizes as correct only those samples containing either one or both of the transmitted protocols, counting a sample as incorrect if classifier outputs any other protocol combination.
TABLE XVII: Global results
TABLE XVIII: Dataset RM_C_O1 (overlap ratio $R=25\%$)
TABLE XIX: Dataset RM_C_O1 (overlap ratio $R=50\%$)
TABLE XX: Dataset RM_C_O2 (overlap ratio $R=25\%$)
Overlapping configuration (incumbent - interferer)
Metrics (%) b - ax b - g b - n n - ax g - ax g - n
Exact Accuracy 69.9 63.8 69.4 62.5 76.9 79.5
Single Exact Accuracy 80.2 76.5 79.4 70.7 82.2 82.6
Single Accuracy 99.8 100.0 99.9 99.9 100.0 100.0
ax detected 80.8 6.4 9.2 92.9 90.6 16.5
b detected 92.9 100.0 96.4 7.0 2.7 0.8
n detected 13.8 17.3 77.7 75.1 15.2 84.5
g detected 6.2 65.7 11.6 22.3 88.9 97.1
noise detected 0.0 0.0 0.0 0.0 0.0 0.0
Overlapping configuration (incumbent - interferer)
Metrics (%) b - ax b - g b - n n - ax g - ax g - n
Exact Accuracy 72.6 97.9 90.9 38.9 84.8 89.1
Single Exact Accuracy 79.6 98.8 94.3 53.1 90.9 94.3
Single Accuracy 99.8 100.0 100.0 99.8 99.9 100.0
ax detected 92.4 0.2 1.5 97.8 99.8 4.0
b detected 84.7 100.0 95.4 19.7 7.6 1.7
n detected 16.5 1.0 96.2 45.9 1.5 98.2
g detected 4.0 98.0 4.2 27.2 86.2 92.7
noise detected 0.0 0.0 0.0 0.1 0.0 0.0
Overlapping configuration (incumbent - interferer)
Metrics (%) b - ax b - g b - n n - ax g - ax g - n
Exact Accuracy 82.8 58.8 70.5 40.3 69.6 96.3
Single Exact Accuracy 87.5 64.0 75.6 51.8 76.3 97.3
Single Accuracy 99.7 100.0 99.9 100.0 100.0 99.9
ax detected 93.1 1.4 5.1 79.8 95.9 1.6
b detected 94.0 100.0 93.3 2.3 1.2 1.1
n detected 12.3 34.7 85.9 74.2 22.5 98.8
g detected 0.3 62.6 19.3 45.9 78.4 98.4
noise detected 0.0 0.0 0.0 0.0 0.0 0.0
Overlapping configuration (incumbent - interferer)
Metrics (%) b - ax b - g b - n n - ax g - ax g - n
Exact Accuracy 49.5 33.4 28.1 97.6 71.4 55.0
Single Exact Accuracy 70.7 58.3 57.1 98.6 76.8 58.7
Single Accuracy 100.0 100.0 99.8 99.9 100.0 100.0
ax detected 53.3 20.8 26.7 99.1 72.6 41.3
b detected 99.7 100.0 99.7 1.2 0.0 0.0
n detected 10.4 21.8 31.3 98.5 23.2 57.7
g detected 19.2 35.0 17.2 0.3 99.7 99.5
noise detected 0.0 0.0 0.0 0.0 0.0 0.0
Overlapping configuration (incumbent - interferer)
Metrics (%) b - ax b - g b - n n - ax g - ax g - n
Exact Accuracy 64.6 49.4 75.6 97.6 81.5 63.2
Single Exact Accuracy 79.5 79.4 84.5 99.1 83.2 65.9
Single Accuracy 99.8 100.0 100.0 100.0 99.9 100.0
ax detected 71.4 9.3 10.1 99.5 86.4 34.1
b detected 96.9 100.0 99.5 0.4 0.0 0.0
n detected 15.2 11.3 83.0 98.2 16.8 68.3
g detected 5.6 51.4 5.3 0.4 98.7 99.7
noise detected 0.0 0.0 0.0 0.0 0.0 0.0
TABLE XVII: Global results
TABLE XVIII: Dataset RM_C_O1 (overlap ratio $R=25\%$)
TABLE XIX: Dataset RM_C_O1 (overlap ratio $R=50\%$)
TABLE XX: Dataset RM_C_O2 (overlap ratio $R=25\%$)
TABLE XXI: Dataset RM_C_O2 (overlap ratio $R=50\%$)
From Table XXI we observe an average of 86.88% correct detection rates for protocol classes that are actually present within the overlapping signals in the test set, while detection rate for protocols not present is on average 7.16%. This demonstrates how our classifier can accurately identify the protocols even in presence of active interference, and it is able to correctly determine both incumbent and interferer transmissions in the majority of cases.
Upon a more detailed analysis of the results, looking at Tables XXI, XXI, XXI and XXI it becomes evident that 802.11b is the protocol most easily identified by the model. This behavior may arise from a number of factors, including distinctiveness of its waveforms compared to others and oversampling applied to signals in this category in order to match other protocols’ configurations in the rest of the dataset.
Another noteworthy observation is the distinct behavior of the model across the two datasets. In dataset RM_C_02, we generally observe superior detection accuracy for the incumbent signal compared to the interferer. However, in dataset RM_C_01, the model deviates from this pattern, exhibiting numerous instances where higher accuracy is achieved for the interferer.
Finally, it is noteworthy to observe the model’s difficulty in identifying the 802.11g protocol particularly when mixed as an interferer with 802.11b. In this specific configuration, the former protocol presents the most significant challenge for the model in terms of identification. Nevertheless, the model overall continues to yield satisfactory results, motivating further research efforts in these overlapping scenarios.